Working alone John can pain a house in 7 hours. If Joe and John paint it together they can finish it in 3 hours.
How long does it take Joe to pain the house?

Answer by stanbon(75887)

Working alone John can paint a house in 7 hours. If Joe and John paint it together they can finish it in 3 hours. h
How long does it take Joe to paint the house?
-----
John rate = 1/7 job/hr
together rate = 1/3 job/hr
---
Joe rate = 1/x job/hr
----------
Equation:
rate + rate = together rate
1/7 + 1/x = 1/3
3x + 21 = 7x
4x = 21
x = 21/4 = 5 1/4 hrs (Joe's time to paint the house alone)
======================
Cheers,
Stan H.